OpenStudy (anonymous):

sum from n=1 -> infinity of (9^n) /(3+ 10^n)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} = (9^n)/ (3 +10^n)\]

OpenStudy (anonymous):

convergent or divergent? I'm not able to use the Satellite's eyeball test for this one

OpenStudy (anonymous):

no but almost ignore the 3

OpenStudy (anonymous):

the n's could be anything

OpenStudy (anonymous):

or since you are doing mathematics and not eyeballing say \(\frac{9^n}{3+10^n}<\frac{9^n}{10^n}\)

OpenStudy (anonymous):

the denominator will always be greater

OpenStudy (anonymous):

and since \(\frac{9^n}{10^n}=.9^n\) you win

OpenStudy (anonymous):

divergent?

OpenStudy (anonymous):

oh no \(.9<1\)

OpenStudy (anonymous):

VICTORY!!!!!

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty}(\frac{9}{10})^n=\frac{.9}{1-.9}=\frac{.9}{.1}=90\]

OpenStudy (anonymous):

oops i mean 9

OpenStudy (anonymous):

it's convergent because it's a geometric series

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes indeed geometric with \(r=.9<1\)

OpenStudy (anonymous):

great thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):